### v1

@@ 3.3.2 Molecular Properties and Molecular Structures
Figure 4: Violin plots of key properties (QED, SA, and Vina docking scores) for 100 ligand molecules generated by MSCoD for each binding pocket of the 7XKJ protein target.
@@ 3.4.1 Effect of the MSCoD Interaction Mechanism on Target-specific Molecule Generation
We observe that the addition of the MSIB module effectively improves the mean Vina Score and median Vina Min values, demonstrating enhanced binding affinity prediction capabilities. Meanwhile, the addition of the MHCA module improves the median Vina Score, showing more stable binding affinity prediction capabilities across different test cases. Furthermore, the combination of these two modules effectively improves QED values, indicating better drug-likeness properties in generated molecules.
@@ 3.4.2 Effect of the Number of Attention Heads in the MHCA on Target-specific Molecule Generation
3.5 3.5 Case Studies
@@
To evaluate the practical applicability of MSCoD, we performed conditional molecular generation experiments on the 7XKJ protein binding pocket, which corresponds to the clinically relevant KRAS G12D mutant. This target is known for its challenging binding environment, including a dynamic, shallow pocket and altered electrostatics due to the G12D mutation(Mao et al.,, 2022; Akkapeddi et al.,, 2023). The 7XKJ structure contains two chemically distinct co-crystallized ligands (GDP and 6IC), providing a robust benchmark for generative model evaluation. We designed two experimental scenarios by defining separate binding pockets around each ligand using a radius-based protocol. For each scenario, MSCoD generated 100 molecules using only the protein structure as input, ensuring unbiased exploration. Performance was compared with MolCRAFT under identical conditions.
@@
MSCoD demonstrates superior performance across all evaluation metrics, generating diverse high-quality molecules suitable for the challenging KRAS G12D binding environment. The result is shown in Figure 4. The co-crystallized ligands establish performance benchmarks: GDP exhibits QED of 0.229, SA of 0.64, and Vina Score of -8.84 kcal/mol, while 6IC achieves QED of 0.319, SA of 0.595, and Vina Score of -10.137 kcal/mol. For the GDP-defined pocket scenario, MSCoD generates molecules with enhanced binding affinity (-9.468 vs -8.84 kcal/mol, 7.1% improvement) and dramatically improved drug-likeness (QED: 0.369 vs 0.229, 61.1% enhancement) compared to the reference GDP ligand. MSCoD also outperforms MolCRAFT with QED scores of 0.369 versus 0.233 (58.5% improvement) and superior Vina Scores of -9.468 versus -9.032. For the 6IC-defined pocket scenario, MSCoD maintains equivalent binding affinity (-10.137 vs -10.137 kcal/mol) while delivering substantial improvements in drug-likeness (QED: 0.559 vs 0.319, 75.2% enhancement) and enhanced synthetic accessibility (SA: 0.68 vs 0.595, 14.3% improvement) compared to the reference 6IC ligand. Additionally, MSCoD outperforms MolCRAFT with QED scores of 0.559 versus 0.422 (32.5% improvement) and comparable binding affinity. These results demonstrate that MSCoD not only matches or exceeds the binding performance of proven co-crystallized ligands but simultaneously addresses their pharmaceutical limitations, with substantial QED improvements (61.1% for GDP, 75.2% for 6IC) indicating significantly more favorable ADMET profiles while maintaining or enhancing binding affinity.
These results demonstrate that MSCoD not only matches or exceeds the binding performance of co-crystallized ligands but also generates molecules with improved pharmaceutical properties, highlighting its potential for challenging oncogenic targets.
@@
In this work, we propose MSCoD, a novel generative framework that significantly advances the modeling of protein-ligand interactions by utilizing multi-scale information bottleneck modules and multi-head cooperative attention mechanism. MSCoD provides a robust and extensible framework for structure-based drug design that enables the precise generation of drug-like molecules with high binding affinity and favorable molecular properties. In experiments, our MSCoD outperforms other state-of-the-art models, and case studies on challenging targets like KRAS G12D further demonstrate its effectiveness. In addition, the general modules developed in this work also hold promise for broader applications, such as protein-protein interaction prediction and other structural biology tasks that require modeling across multiple spatial and temporal scales. In summary, this work will expedite the drug discovery process.

### v2

@@ 3.3.1 Target Binding Affinity
Figure 4: Bond length distribution of reference and generated molecules by autoregressive models (upper row) and nonautoregressive models (lower row) for top-5 frequent bond types
(C-C-C Bond Angle)
(C:C:N Bond Angle)
(C-C-O Bond Angle)
Figure 5: Bond angle distributions of generated molecules compared with reference molecules.
(C-C-C-C Torsion Angle)
(C:C:C:C Torsion Angle)
Figure 6: Torsion angle distributions of generated molecules compared with reference molecules.
@@ 3.3.2 Molecular Properties and Molecular Structures
Additionally, we also analyzed the distributions of bond lengths, bond angles, and torsion angles of the generated ligands to assess structural plausibility. Figure 4 compares bond-length distributions for the five most frequent bond types; the distributions of MSCoD-generated molecules closely match those of the reference set across all bond classes, indicating accurate control of interatomic distances. Figures 5 and 6 present bond-angle and torsion-angle distributions, respectively; the generated molecules show strong agreement with reference distributions for these geometric descriptors, further validating the model’s ability to reproduce realistic 3D geometry. Together, these results demonstrate that MSCoD not only improves binding affinity and drug-like properties but also generates chemically and geometrically plausible molecular structures, supporting its suitability for practical structure-based drug design.
@@ 3.4.1 Effect of the MSCoD Interaction Mechanism on Target-specific Molecule Generation
We observe that adding the MSIB module improves the mean Vina Score and median Vina Min values, demonstrating enhanced binding affinity prediction. Meanwhile, adding the MHCA module improves the median Vina Score, indicating more stable binding affinity prediction across different test cases. Furthermore, the combination of these two modules effectively improves QED values, indicating better drug-likeness properties in generated molecules.
@@ 3.4.2 Effect of the Number of Attention Heads in the MHCA on Target-specific Molecule Generation
Figure 7: Property distributions (QED, SA, and Vina Score) for the 100 molecules generated by MSCoD for the two distinct binding pockets of the KRAS G12D (7XKJ) target, defined by the co-crystallized ligands GDP and 6IC, respectively.
[…]
GIN (MSCoD) 0.492 0.242 0.837 0.698 0.888 0.402 0.162 0.875 0.873 0.888
[…]
GAT (MSCoD) 0.500 0.250 0.831 0.673 0.874 0.425 0.181 0.857 0.851 0.873
[…]
GCN (MSCoD) 0.512 0.262 0.821 0.672 0.874 0.395 0.156 0.879 0.875 0.891
[…]
GAT_GCN (MSCoD) 0.508 0.258 0.827 0.695 0.887 0.387 0.150 0.884 0.878 0.892
[…]
Table 4: Drug-target affinity (DTA) prediction performance on the Davis and Kiba datasets. The methods are GraphDTA variants, where ”(MSCoD)” indicates the baseline has been augmented with the MSCoD block (MSIB and MHCA). Boldface highlights the superior value in each baseline/augmented pair.
3.5 3.5 Case Study: Generation for the Challenging 7XKJ Target
@@
To evaluate the practical applicability of MSCoD, we performed conditional molecular generation experiments on the 7XKJ protein binding pocket, which corresponds to the clinically relevant KRAS G12D mutant. This target is known for its challenging binding environment, including a dynamic, shallow pocket and altered electrostatics due to the G12D mutation(mao2022kras; akkapeddi2023exploring). The 7XKJ structure contains two chemically distinct co-crystallized ligands (GDP and 6IC), providing a robust benchmark for generative model evaluation. We designed two experimental scenarios by defining separate binding pockets around each ligand using a radius-based protocol. For each scenario, MSCoD generated 100 molecules using only the protein structure as input, ensuring unbiased exploration. Performance was compared with MolCRAFT under identical conditions.
@@
MSCoD demonstrated superior performance across all evaluation metrics, generating diverse, high-quality molecules suitable for the challenging KRAS G12D binding environment. The co-crystallized ligands establish performance benchmarks: GDP exhibits a Vina Score of -8.84 kcal/mol and QED of 0.229, while 6IC achieves -10.137 kcal/mol and 0.319, respectively. In the GDP-defined pocket scenario, MSCoD generated molecules with both enhanced binding affinity (-9.468 vs. -8.84 kcal/mol, a 7.1% improvement) and significantly improved drug-likeness (QED: 0.369 vs. 0.229, a 61.1% enhancement) compared to the reference GDP ligand. It also outperformed MolCRAFT with stronger Vina Scores (-9.468 vs. -9.032) and higher QED (0.369 vs. 0.233). In the 6IC-defined pocket scenario, MSCoD maintained an equivalent top-tier binding affinity (-10.137 kcal/mol) while delivering substantial improvements in drug-likeness (QED: 0.559 vs. 0.319, a 75.2% enhancement) and outperforming MolCRAFT (QED: 0.559 vs. 0.422). These results demonstrate that MSCoD not only matches or exceeds the binding performance of proven co-crystallized ligands but also generates molecules with superior pharmaceutical properties, highlighting its potential for challenging oncogenic targets.
3.6 3.6 Transferability Study: Enhancing DTA Prediction with MSIB and MHCA
To evaluate the transferability of our proposed components, we integrated the MSIB and MHCA modules as a single, combined block—hereafter referred to as the MSCoD block—into four GraphDTA variants (GIN, GAT, GCN, and GAT_GCN) for drug–target affinity (DTA) prediction (nguyen2021graphdta). The experiments were strictly controlled to isolate the effect of this block: baseline architectures and all original hyperparameters remained unchanged. All models were trained under an identical protocol on the Davis and Kiba datasets, with results averaged over three independent runs (Table 3.4.2).
To assess performance on the DTA task, we employ a set of standard regression and ranking metrics. The Root Mean Square Error (RMSE) and Mean Squared Error (MSE) quantify prediction accuracy by measuring the deviation between predicted and actual binding affinity values; lower values indicate a more accurate model. For evaluating correlation, the Pearson coefficient measures the linear relationship, while the Spearman’s rank coefficient assesses the monotonic relationship, indicating how well the model’s predictions rank the compounds. Finally, the Concordance Index (CI) directly evaluates the probability that the model correctly ranks any two randomly chosen compounds. In drug discovery, strong performance in ranking metrics like Spearman and CI is particularly crucial, as it reflects a model’s practical utility in correctly prioritizing potent compounds from a large candidate pool.
Across all encoder variants, augmenting GraphDTA with the MSCoD block led to consistent and measurable improvements. Notably, these gains were achieved without any further hyperparameter tuning, highlighting the ”plug-and-play” utility of our approach. For example, the GAT_GCN backbone benefited substantially: on Davis, RMSE decreased from 0.532 to 0.508 (a 4.5% reduction) while CI increased from 0.874 to 0.887; on Kiba, RMSE decreased from 0.408 to 0.387 (a 5.1% reduction) and CI rose from 0.873 to 0.892. Comparable gains were observed for the GIN, GAT, and GCN backbones, indicating that the performance enhancement is not specific to a single encoder design.
The magnitude of the gains is modest but consistent (typical RMSE reductions of  0.02–0.05 and small increases in Pearson/Spearman/CI), which suggests that MSIB improves hierarchical feature extraction while MHCA strengthens asymmetric protein→ligand interaction modeling. Under an identical training protocol, these results support the modules’ practical transferability: inserting MSIB and adding MHCA to diverse GraphDTA backbones enhances affinity-prediction accuracy and ranking, demonstrating the utility of MSCoD components for cross-task applications in structure-based drug discovery.
@@
In this work, we propose MSCoD, a novel generative framework that significantly advances the modeling of protein-ligand interactions by utilizing multi-scale information bottleneck modules and multi-head cooperative attention mechanism. MSCoD provides a robust and extensible framework for structure-based drug design that enables the precise generation of drug-like molecules with high binding affinity and favorable molecular properties. Experimental studies show that MSCoD achieves superior overall performance on structure-based molecular generation benchmarks. Case studies on challenging targets such as KRAS G12D further demonstrate its effectiveness for difficult protein targets, and the MSIB and MHCA modules also transfer effectively to drug–target affinity prediction, enhancing predictive performance on standard datasets. In summary, this work has the potential to expedite the drug discovery process.